OpenStudy (anonymous):

explain this quadratic equation in standard form. x squared -10=3x? please show work :)

OpenStudy (anonymous):

The equation is x^2-2x+1=0

OpenStudy (callisto):

both sides -3x so that the right hand side becomes 0

OpenStudy (callisto):

x squared -10=3x x squared -10 -3x=3x-3x x squared -3x-10=0

OpenStudy (callisto):

@kumar2006 don't give the wrong answer without explanations

OpenStudy (anonymous):

First of all that wasn't the answer, it was an equation

OpenStudy (callisto):

then don't give the inappropriate response or you should explain it if it is related to the question

OpenStudy (anonymous):

Look man I don't got time to argue with you online and I will do what I want. So who the hell are you to tell me what to do or not? I don't care what you really say man because I know what I am doing. you need to stop being a online thug.

OpenStudy (anonymous):

I got up to level 44 and have 484 medals for a reason. Think?